# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v1

@@ 5.5 Limitations
While NGDINO addresses the multi-target and no-target challenges, several challenges remain. Figure 8 shows typical NGDINO failure cases, highlighting challenges in the RefDrone dataset. The examples show rich contextual expressions, challenging backgrounds, and small-scale object detection. These challenging cases reflect real-world applications and highlight the areas for future improvement.
@@ 6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes. The dataset is constructed using the proposed RDAgent, a semi-auto annotation framework leveraging a multi-agent system. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in the RefDrone dataset. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to expand the benchmark to include referring expression segmentation and referring expression tracking tasks.
Supplementary Material
We provide supplementary material for further analysis, organized as follows:
@@
    Color attribute analysis of RefDrone. (Appendix A)
@@
    Details of baseline method. (Appendix B)
@@
    Additional dataset examples. (Appendix C)
@@
    Resutlts w.r.t different object scales. (Appendix D)
@@
    Prompts and examples in RDAgent. (Appendix E)
Appendix A Dataset color attribute analysis
In the RefDrone dataset, color attributes are present in 69% expressions. We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow. During the annotation process, RDAgent expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver. The distribution of these color terms across the expressions is illustrated in Figure 9.
Figure 9: Distribution of color terms in RefDrone expressions.
Appendix B Details of baseline methods
@@
      pretrained on Object365, GoldG, GRIT, V3Det.
@@
      pretrained on Object365, GoldG, V3Det.
@@
    • SPHINX-v2, Griffon: 13B parameters.
Appendix C Dataset examples
@@
Appendix D The results of different object scales.
We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) of RefDrone dataset in Table 8.
Table 8: Fine-tuned results w.r.t. different object scales.
Methods Accs Accm Accl Methods Accs Accm Accl GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99 GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18 NGDINO-T 10.84 23.38 40.13 RDAgent 32.54 47.97 47.17
Appendix E Prompts and examples for RDAgent
In this section, we provide the prompts and examples employed in RDAgent. Table 9 presents the prompt construction process for expression generation (Step 3), which includes the system prompt and few-shot in-context learning examples. One in-context learning example is illustrated in Table 10. The system prompts used for each step are detailed in Table 11. Additionally, the system prompts for the feedback mechanism are presented in Table 12.
Figure 10: Dataset samples from RefDrone.
Table 9: Illustration of RDAgent’s prompt construction for expression generation (Step 3). Few-shot in-context-learning examples are from fewshot_samples. A representative example is shown in Figure 10
messages = [ {"role":"system", "content": f"""As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Your task is to classify the provided objects based on various characteristics, while also substantiating your classification. This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects. To refer to a specific object, use the provided coordinates directly. Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions."""} ] for  sample in fewshot_samples:
       messages.append({"role":"user", "content":sample[‘context’]})
messages.append({"role":"assistant", "content":sample[‘response’]} )
@@ Appendix E Prompts and examples for RDAgent
Table 10: One example to illustrate the input and output in RDAgent Step 3 (expression generation). The top block shows the input contexts such as image, object, caption and color attributed. The bottom block shows the generated annotations.
@@
Table 11: System prompts employed at different stages of RDAgent, replacing the default system messages shown in Table 9
@@
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation
@@
Your task is to confirm whether the description exclusively relates to the described objects without including any others in the visual. Respond ”yes” if it matches, or ”no” with an explanation if it does not.
Table 12: RDAgent system prompts for feedback mechanism. Differences from Table 11 are highlighted
@@
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation with feedback mechanism

### v2

@@ 5.3 Limitations
While NGDINO addresses the multi-target and no-target challenges, several challenges remain. Figure 7 shows some typical NGDINO failure cases, highlighting challenges in RefDrone. The examples show rich contextual reasoning, challenging backgrounds, and small-scale object detection. These challenging cases reflect real applications and highlight areas for future improvement.
@@ 6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes. The dataset is constructed using the RDAnnotator framework, an efficient semi-automated annotation pipeline combining LMM-based modules and human oversight. RDAnnotator provides high-quality annotations while reducing human effort. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in RefDrone. Experimental results demonstrate consistently low performance across existing approaches, underscoring the inherent difficulty of the RefDrone benchmark. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to extend RefDrone to larger-scale scenarios and introduce additional tasks such as referring expression segmentation and referring expression tracking. We believe RefDrone will serve as a valuable benchmark for advancing research in drone-based REC tasks.
Appendix A Appendix
We provide the following appendices for further analysis:
@@ Appendix A Appendix
    Details of the baseline method. (Appendix A.1)
@@
    Implementation details. (Appendix A.2)
@@
    Color attribute analysis of RefDrone. (Appendix A.3)
@@
    Additional dataset examples. (Appendix A.4)
@@
    Results with respect to different object scales. (Appendix A.5)
  6. 6.
    Results on RefCOCO/+/g datasets. (Appendix A.6)
  7. 7.
    Visualization results analysis. (Appendix A.7)
  8. 8.
    Prompts and examples in RDAnnotator. (Appendix A.8)
A.1 Details of baseline methods
@@
      pretrained on Objects365, GoldG, GRIT, V3Det.
@@
      pretrained on Objects365, GoldG, V3Det.
@@
    • SPHINX-v2: 13B parameters.
    • Griffon: 13B parameters.
A.2 Implementation details
NGDINO implementation. The NGDINO adopts a two-stage procedure to maintain training stability. First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [34] parameters. We then fine-tune the entire model. We also evaluate it on the gRefCOCO [28] dataset, which includes no-target and limited multi-target samples. Additionally, we evaluate it on RSVG [56], an aerial-view dataset with multi-target challenges.
Zero-shot evaluation details. For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers. The implementations and checkpoints for GLIP and GDINO are obtained from MMDetection [6].
Fine-tuning evaluation details. Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison. We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples. For LMMs, we employ the LoRA [17] fine-tuning strategy and follow the instruction tuning data structures. All fine-tuning experiments are run for 5 epochs using 4 NVIDIA A100 GPUs.
A.3 Details of color categorization.
We design a hybrid color extraction pipeline consisting of two main components: (1) a WideResNet-101 classifier [55] trained on HSV color space-based labels. The color labels are refined via manual validation, ensuring robustness to annotation noise and label ambiguity. (2) An LMM-based color consistency module that validates color predictions through structured prompts. The latter cross-checks the classifier’s outputs to mitigate ambiguities caused by illumination variance or partial occlusion.
In the RefDrone dataset, color attributes are present in 69% of expressions. We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow. During the annotation process, RDAnnotator expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver. The distribution of these color terms across the expressions is illustrated in Figure 8.
Figure 8: Distribution of color terms in RefDrone expressions.
A.4 Dataset examples
@@
A.5 The results of different object scales.
We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) for the RefDrone dataset in Table 9.
Table 9: Fine-tuned results with respect to different object scales.
Methods Accs Accm Accl Methods Accs Accm Accl
GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99
GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18
NGDINO-T 10.84 23.38 40.13 RDAnnotator 32.54 47.97 47.17
A.6 Results on RefCOCO/+/g datasets.
Since the RefCOCO, RefCOCO+, RefCOCOg datasets contain only one instance per expression, the proposed NGDINO leverages the number branch primarily to address multi-instance and no-instance scenarios. As a result, the performance of NGDINO is relatively similar to that of GDINO on these datasets.
Table 10: Results on RefCOCO/+/g datasets.
RefCOCO RefCOCO+ RefCOCOg
TestA TestB TestA TestB Val Test
MDETR 90.4 82.67 85.52 72.96 83.35 83.31
GDINO-T 91.4 86.6 87.5 74.0 85.5 85.8
NGDINO-T 91.5 86.5 87.8 74.7 85.3 85.8
A.7 Visualization results compared NGDINO with baseline GDINO.
Figure 9 shows visualization results comparing NGDINO and the baseline GDINO on the proposed RefDrone dataset, demonstrating NGDINO’s effectiveness with multi-target samples.
Figure 9: Visualization results comparing (a) GDINO [34] and (b) NGDINO, where red boxes indicate true positives and yellow boxes denote false negatives.
A.8 Prompts and examples for RDAnnotator
In this section, we provide the prompts and examples employed in RDAnnotator. Table 11 presents the prompt construction process for expression generation (Step 3), which includes the system prompt and few-shot in-context learning examples. One in-context learning example is illustrated in Table 12. The system prompts used for each step are detailed in Table 13. Additionally, the system prompts for the feedback mechanism are presented in Table 14.
Figure 10: Dataset examples from RefDrone.
Table 11: Illustration of RDAnnotator’s prompt construction for expression generation (Step 3). Few-shot in-context-learning examples are from fewshot_samples. A representative example is shown in Table 12
@@ A.8 Prompts and examples for RDAnnotator
Table 12: One example to illustrate the input and output in RDAnnotator Step 3 (expression generation). The top block shows the input contexts such as image, object, caption and color attributed. The bottom block shows the generated annotations.
@@
Table 13: System prompts employed at different stages of RDAnnotator, replacing the default system messages shown in Table 11
@@
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with "Yes" when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise "No." System prompt: Step 3 expression generation
@@
Your task is to confirm whether the description exclusively relates to the described objects without including any others in the visual. Respond "yes" if it matches, or "no" with an explanation if it does not.
Table 14: RDAnnotator system prompts for feedback mechanism. Differences from Table 13 are highlighted
@@
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with "Yes" when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise "No." System prompt: Step 3 expression generation with feedback mechanism